OpenStudy (anonymous):

Silicon tetrachloride is used in the electronics industry to make elemental silicon for computer chips. Silicon tetrachloride is prepared from silicon dioxide, graphitic carbon, and chlorine gas: SiO2(s)+2C(s)+2Cl2(g)−−\>SiCl4(l)+2CO(g) If the reaction occurs with a 90.0% yield, what mass of silicon tetrachloride can be prepared from 51.9 g of each starting material? Enter units. Im new to chemistry and this site a friend told me i should do this since i needed help.. PLEASE HELPPP

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20100602162644AAl9t3v same question is asked there with different value . May this help u .

OpenStudy (aaronq):

Convert 51.9 g to moles, determine which reactant you have less of (i.e. find the limiting reactant). use those moles to build a ratio, and solve for the moles of product. ------------------------------------------------------------- e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.